# Functional Regression with Nonstationarity and Error Contamination: Application to the Economic Impact of Climate Change

### v1

@@ 3 Proposed model
Under Assumption 1E, the above is a well defined bounded linear operator acting on $\mathcal{H}$ (see Section 2.3. of Beare et al.,, 2017). We hereafter let $\mathfrak{F}_{t}$ be the filtration given by
$\mathfrak{F}_{t}=\sigma(\{u_{s}\}_{s\leq t-1},\{u_{s}^{N}\}_{s\leq t},\{u_{s}^{S}\}_{s\leq t}).$ (5)
@@ 4.1 Autocovariance-based FPCA
We then can define its inverse on the restricted domain $\operatorname{ran}(\sum_{j=1}^{{\mathsf{K}}}\bm{{\Pi}}_{j}[\widehat{D}_{\kappa}])$ for ${\mathsf{K}}>0$ as follows:
@@
In the case where $\kappa=0$ (and thus $\widehat{C}_{\kappa}(\widehat{D}_{\kappa})^{-1}_{\mathsf{K}}=\sum_{j=1}^{\mathsf{K}}\bm{\mathsf{\lambda}}^{-1}_{j}[\widehat{C}_{\kappa}]\bm{{\Pi}}_{j}[\widehat{C}_{\kappa}]$) and $z_{t}=y_{t}$, (7) becomes identical to the standard FPCA-based estimator considered in the literature concerning stationary functional time series (see e.g., Park and Qian,, 2012). However, as may be deduced from the results of Seong and Seo, (2025), this estimator is subject to the presence of measurement errors and may not be a consistent estimator of $f$. In our context, $f$ on the subspace $\mathcal{H}^{S}=\operatorname{ran}{\rm P}^{S}$ is not generally consistently estimated, which results from the fact that the sample covariance ${\rm P}^{S}\widehat{C}_{0}{\rm P}^{S}$ suffers from non-negligible contamination by measurement errors (note that ${\rm P}^{S}\widehat{C}_{0}{\rm P}^{S}$ contains the component $T^{-1}\sum_{t=1}^{T}{\rm P}^{S}e_{t}\otimes{\rm P}^{S}e_{t}$, which is non-negligible) and thus is an inconsistent estimator of its true counterpart (i.e., $\mathbb{E}[{\rm P}^{S}x_{t}\otimes{\rm P}^{S}x_{t}]$). On the other hand, under Assumption E(a), we may deduce that ${\rm P}^{S}\widehat{C}_{\kappa}{\rm P}^{S}$ for $\kappa\geq 1$ does not suffer from such a serious contamination. This is the reason why we will mainly consider $\kappa\geq 1$ and construct our proposed estimator using the sample autocovariance operator. Computation of $(\widehat{D}_{\kappa})^{-1}_{\mathsf{K}}$ requires determining the number of retained eigenvectors, $\mathsf{K}$}. Subsequently, we will require $\mathsf{K}$} to grow without bound depending on certain sample eigenvalues, but for now we assume only the following, required for the first few main results:
@@ 4.2 Proposed estimator
where $\mathfrak{Y}_{T}=O_{p}(1)$ (see Remark 4.3 for a detailed expression of $\mathfrak{Y}_{T}$), $V_{1}=_{d}\int W^{N}\otimes W^{N}$ and $V_{2}=_{d}\int W^{N}\otimes dW^{u}$. If there is no measurement error (i.e., $e_{t}=0$), the $\mathfrak{Y}_{T}=0$.
Theorem 4.2 demonstrates that the proposed estimator $\widehat{f^{N}_{\kappa}}$ is a consistent estimator of $f^{N}$, and the asymptotic bias is of order $T^{-1}$; this result parallels that of the standard least squares-type estimator for the cointegrating relationship in the finite-dimensional case. We present remarks that contain some complementary results to Theorem 4.2.
@@
In fact, our proof of Theorem 4.2 given in Section D accommodates this more general case.
We next study the asymptotic properties of $\widehat{f^{S}_{\kappa}}$ as an estimator of $f^{S}(=f{\rm P}^{S})$. As in the standard FPCA-based estimators, our proposed estimator given in (14) is defined on a finite dimensional eigenspace of $\widehat{D}_{\kappa}^{S}$. Using the result that $\widehat{{\rm P}}^{S}_{\kappa}-{\rm P}^{S}=O_{p}(T^{-1})$ (see Theorem 4.1), we may deduce that $\widehat{D}_{\kappa}^{S}$ is a consistent estimator of $D_{\kappa}^{S}$, defined below:
@@
Obviously, (21) may be used for inference on $f\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}(\zeta)$, where $\widehat{{\rm P}}^{K}_{\kappa}(\zeta)$ is naturally understood as the optimal approximation of $\zeta$ using the eigenvectors of $\widehat{D}_{\kappa}$. For example, when $\mathcal{H}_{y}=\mathcal{H}$ (and hence $y_{t}$ is function-valued), we may construct the 95% confidence interval of $\langle f\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}(\zeta),\varphi\rangle$ for some $\varphi\in\mathcal{H}_{y}$ using the asymptotic normality result (21), as follows::
$\langle\widehat{f}_{\kappa}\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}(\zeta),\varphi\rangle\pm 1.96\sqrt{{\theta}_{{\mathsf{K}}_{S}}\langle\widetilde{C}_{u}\varphi,\varphi\rangle/T},$
where the unknown quantities ${\theta}_{{\mathsf{K}}_{S}}$ and $\widetilde{C}_{u}$ can be replaced by reasonable estimators that can be easily computed from our proposed estimator $\widehat{f}_{\kappa}$ without affecting asymptotic validity (see Corollary C.1 of the Appendix). However, practitioners may want to avoid being interfered by a random projection $\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}$ and implement a direct statistic inference on $\langle f(\zeta),\varphi\rangle$, rather than on $\langle f\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}(\zeta),\varphi\rangle$. In fact, we will show that, under additional assumptions (requring “smoothness” of $f$ and $\zeta$), the asymptotic normality result (21) still holds even when $f\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}$ is replaced by $f$. Based on this result, we can implement statistical inference without the influence of the random projection $\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}$. This will be more detailed in the next section.
@@ 4.3 Statistical inference: local confidence bands of a partial effect
The above theorem requires $\theta_{{\mathsf{K}}_{S}}(\zeta)\to\infty$. This is likely to be true for many possible choices of $\zeta$; for example, if $\zeta$ is arbitrarily chosen from $\mathcal{H}$, $\mathbb{P}\{\theta_{{\mathsf{K}}_{S}}(\zeta)<c<\infty\}\to 0$ as $\mathsf{K}$$\to\infty$ since $\theta_{{\mathsf{K}}_{S}}(\zeta)$ is only convergent on a strict subspace of $\mathcal{H}$. A more detailed discussion on this result can be found in e.g., Remark 4 of Seong and Seo, (2025).
Even if all the assumptions required for Theorem 4.4 hold, $\widehat{f}_{\kappa}-f$ converges to a Gaussian random element at a rate depending on $\zeta$ and thus it is not generally possible to construct a uniform confidence band of $f$ from Theorem 4.4 (see Remark 4.5). However, it may be possible to construct a local (or locally approximate) confidence band, which is naturally interpreted. We first note that $f(\zeta)$ may be understood as a partial effect on $y_{t}$ of a perturbation $\zeta$ in $x_{t}$, which is often of interest in practice. If $\mathcal{H}_{y}=\mathbb{R}$ and hence $y_{t}$ is real-valued, $f(\zeta)$ is a real-valued effect on $y_{t}$ of a perturbation $\zeta$, and in this case we may directly use (22) for statistical inference by replacing $\widetilde{C}_{u}$ and $\theta_{{\mathsf{K}}_{S}}$ with their sample counterparts (see Corollary C.1 of the Appendix). Now suppose that $\mathcal{H}_{y}=\mathcal{H}$. In this case, $f(\zeta)$ is a function defined on $a_{1},a_{2}]$, and we may construct a sequence of confidence intervals for local averages of $f(\zeta)$. Specifically, let $\mathcal{I}_{j}=(b_{j+1}-b_{j})^{-1}1\{u\in[b_{j},b_{j+1}]\}$ for some $b_{j}$ and $b_{j+1}$ with $a_{1}\leq b_{j}<b_{j+1}\leq a_{2}$. Then $\langle f(\zeta),\mathcal{I}_{j}\rangle=(b_{j+1}-b_{j})^{-1}\int_{b_{j}}^{b_{j+1}}f(\zeta)(s)ds$ computes the local average of $f(\zeta)$ on the interval $b_{j},b_{j+1}]$. Using the results given in Theorem 4.4, we know that
$\sqrt{T/\theta_{{\mathsf{K}}_{S}}(\mathcal{I}_{j})}(\langle(\widehat{f}_{\kappa}(\zeta)-f(\zeta)),\mathcal{I}_{j}\rangle\to_{d}N(0,\langle\mathcal{I}_{j},\widetilde{C}_{u}\mathcal{I}_{j}\rangle).$ (23)
Note that $\langle\mathcal{I}_{j},\widetilde{C}_{u}\mathcal{I}_{j}\rangle$ and $\theta_{{\mathsf{K}}_{S}}$ can also be replaced by their sample counterparts (Corollary C.1 of the Appendix), allowing construction of an asymptotically valid confidence band for the local average using (23). This can be applied to overlapping or non-overlapping sequences of intervals $\{\mathcal{I}_{j}\}_{j=1}^{M}$ with $\cup_{j=1}^{M}\mathcal{I}_{j}=[a_{1},a_{2}]$. These confidence intervals are readily interpretable and can be constructed even if $\theta_{{\mathsf{K}}_{S}}(\mathcal{I}_{j})$ diverge at different rates across $j$.
@@ 4.4 The model with an intercept
In the previous sections, we developed statistical inferential methods for the case where $\mathbb{E}[y_{t}]=\mathbb{E}[x_{t}]=0$ for simplicity. However, in practice, a nonstationary time series may include a nonzero intercept or drift, and hence our observations may be given by $\{\mu_{y}+y_{t}\}_{t\geq 1}$ and $\{\mu_{x}+\tilde{x}_{t}\}_{t\geq 1}$ for some unknown $\mu_{y}$ and $\mu_{x}$. To accommodate this scenario, one may consider the model with a deterministic term as follows: for $\mu\in\mathcal{H}_{y}$,
$y_{t}=\mu+f(x_{t})+u_{t},$ (24)
where $\mu=f(\mu_{x})-\mu_{y}\in\mathcal{H}$. With a straightforward modification, we can still achieve consistent estimation of $f$ and extend the statistical inference on $f(\zeta)$ given in Section 4.3. More specifically, inference for this case can be implemented using the centered (demeaned) variables ${y}_{c,t}=y_{t}-\bar{y}_{T}$ and $\tilde{x}_{c,t}=x_{t}+\eta_{t}-\bar{x}_{T}-\bar{\eta}_{T}$, where $\bar{y}_{T}=T^{-1}\sum_{t=1}^{T}y_{t}$, and $\bar{x}_{T}$ and $\bar{\eta}_{T}$ are similarly defined. The proposed estimator is given as follows: $\widehat{f}_{c,\kappa}=\widehat{f}_{c,\kappa}^{N}+\widehat{f}_{c,\kappa}^{S},$ where
@@
where $\theta_{c,{\mathsf{K}}_{S}}$ is defined similarly to $\theta_{{\mathsf{K}}_{S}}$ in (D.3), but with $C_{\kappa}^{S}$ and the other operators ($\widetilde{C}_{0}^{S}$, $D_{\kappa}^{S}$, $(D_{\kappa}^{S})_{{\mathsf{K}}_{S}}^{-1}$) depending on $C_{\kappa}^{S}$ being computed from ${\rm P}^{S}\tilde{x}_{t-k}-\mathbb{E}[{\rm P}^{S}\tilde{x}_{t-k}]$ instead of ${\rm P}^{S}\tilde{x}_{t-k}$. Of course, as in the previous case, $\theta_{c,{\mathsf{K}}_{S}}$ and $\widetilde{C}_{u}$ can be replaced by their sample counterparts without affecting the asymptotic result given by (25), and thus we may implement statistical inference on $f(\zeta)$ in practice. A more detailed discussion including theoretical justification of these results is given in Section C.1.2 of the Appendix.
@@ 5.1 Monte Carlo simulation
We implemented simulation experiments to compare the autocovariance-based estimator with the standard covariance-based one in the presence of measurement errors. We, however, postpone detailing them to Section B of the Appendix to focus more on the real-data analysis for studying the economic impact of climate change, which more effectively illustrates the empirical relevance and usefulness of our proposed methods.

### v3

@@ 3 Proposed model
Under Assumption 1E, the above is a well-defined bounded linear operator acting on $\mathcal{H}$ (see Section 2.3. of Beare et al.,, 2017). We hereafter let $\mathfrak{F}_{t}$ be the filtration given by
$\mathfrak{F}_{t}=\sigma(\{u_{s}\}_{s\leq t},\{u_{s}^{N}\}_{s\leq t+1},\{u_{s}^{S}\}_{s\leq t+1}).$ (4)
@@ 4.1 Autocovariance-based FPCA
We then define its inverse on the restricted domain $\operatorname{ran}(\sum_{j=1}^{{\mathsf{K}}}\bm{{\Pi}}_{j}[\widehat{D}_{\kappa}])$ for ${\mathsf{K}}>0$ as follows:
@@
In the case where $\kappa=0$ (and thus $\widehat{C}_{\kappa}(\widehat{D}_{\kappa})^{-1}_{\mathsf{K}}=\sum_{j=1}^{\mathsf{K}}\bm{\mathsf{\lambda}}^{-1}_{j}[\widehat{C}_{0}]\bm{{\Pi}}_{j}[\widehat{C}_{0}]$) and $z_{t}=y_{t}$, (6) becomes identical to the standard FPCA-based estimator considered in the literature concerning stationary functional time series (see e.g., Park and Qian,, 2012). However, as the results of Seong and Seo, (2025) suggest, this estimator is affected by measurement errors and may not be a consistent estimator of $f$. In our context, $f$ on the subspace $\mathcal{H}^{S}=\operatorname{ran}{\rm P}^{S}$ is not generally consistently estimated, which results from the fact that the sample covariance ${\rm P}^{S}\widehat{C}_{0}{\rm P}^{S}$ suffers from non-negligible contamination by measurement errors (note that ${\rm P}^{S}\widehat{C}_{0}{\rm P}^{S}$ contains the component $T^{-1}\sum_{t=1}^{T}{\rm P}^{S}e_{t}\otimes{\rm P}^{S}e_{t}$, which is non-negligible) and thus is an inconsistent estimator of its true counterpart (i.e., $\mathbb{E}[{\rm P}^{S}x_{t}\otimes{\rm P}^{S}x_{t}]$). On the other hand, under Assumption E(a), we may deduce that ${\rm P}^{S}\widehat{C}_{\kappa}{\rm P}^{S}$ for $\kappa\geq 1$ does not suffer from such serious contamination. This is why we will mainly consider the case $\kappa\geq 1$ and construct our proposed estimator using the sample autocovariance operator. Computing $(\widehat{D}_{\kappa})^{-1}_{\mathsf{K}}$ requires determining the number of retained eigenvectors, $\mathsf{K}$}. Subsequently, we will require $\mathsf{K}$} to grow without bound depending on certain sample eigenvalues, but for now we assume only the following, required for the first few main results:
@@ 4.2 Proposed estimator
where $\mathfrak{Y}_{T}=O_{p}(1)$ (see Remark 4.3 for a detailed expression of $\mathfrak{Y}_{T}$), $V_{1}=_{d}\int W^{N}\otimes W^{N}$ and $V_{2}=_{d}\int W^{N}\otimes dW^{u}$. If there is no measurement error (i.e., $e_{t}=0$), then $\mathfrak{Y}_{T}=0$.
Theorem 4.2 demonstrates that the proposed estimator $\widehat{f^{N}_{\kappa}}$ is a consistent estimator of $f^{N}$, and the asymptotic bias is of order $T^{-1}$; this result parallels that of the standard least squares-type estimator for the cointegrating relationship in the finite-dimensional case. We present remarks containing complementary results to Theorem 4.2.
@@
In fact, our proof of Theorem 4.2 given in Section D of the Supplementary Material accommodates this more general case.
We next study the asymptotic properties of $\widehat{f^{S}_{\kappa}}$ as an estimator of $f^{S}(=f{\rm P}^{S})$. As with standard FPCA-based estimators, our proposed estimator given in (13) is defined on a finite-dimensional eigenspace of $\widehat{D}_{\kappa}^{S}$. Using the result that $\widehat{{\rm P}}^{S}_{\kappa}-{\rm P}^{S}=O_{p}(T^{-1})$ (see Theorem 4.1), we may deduce that $\widehat{D}_{\kappa}^{S}$ is a consistent estimator of $D_{\kappa}^{S}$, defined below:
@@
Obviously, (20) may be used for inference on $f\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}(\zeta)$, where $\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}(\zeta)$ is naturally understood as the optimal approximation of $\zeta$ using the eigenvectors of $\widehat{D}_{\kappa}$. For example, when $\mathcal{H}_{y}=\mathcal{H}$ (and hence $y_{t}$ is function-valued), we may construct the 95% confidence interval of $\langle f\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}(\zeta),\varphi\rangle$ for some $\varphi\in\mathcal{H}_{y}$ using the asymptotic normality result (20), as follows::
$\langle\widehat{f}_{\kappa}\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}(\zeta),\varphi\rangle\pm 1.96\sqrt{{\theta}_{{\mathsf{K}}_{S}}\langle\widetilde{C}_{u}\varphi,\varphi\rangle/T},$ (21)
where the unknown quantities ${\theta}_{{\mathsf{K}}_{S}}$ and $\widetilde{C}_{u}$ can be replaced by reasonable estimators that can be easily computed from our proposed estimator $\widehat{f}_{\kappa}$ without affecting asymptotic validity (see Corollary C.1 of the Supplementary Material). However, practitioners may want to avoid being interfered by a random projection $\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}$ and implement a direct statistic inference on $\langle f(\zeta),\varphi\rangle$ rather than on $\langle f\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}(\zeta),\varphi\rangle$. We will show that, under additional assumptions (requring “smoothness” of $f$ and $\zeta$), the asymptotic normality result (20) still holds even when $f\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}$ is replaced by $f$. This allows statistical inference without the influence of the random projection $\widehat{{\rm P}}^{{\mathsf{K}}}_{\kappa}$, as discussed in more detail in the next section.
@@ 4.3 Statistical inference: local confidence bands of a partial effect
The above theorem requires that $\theta_{{\mathsf{K}}_{S}}(\zeta)\to\infty$. This is likely to be true for many possible choices of $\zeta$. As discussed in the literature (see, e.g., Remark 4 of Seong and Seo,, 2025, and references therein), $\theta_{{\mathsf{K}}_{S}}(\zeta)$ is convergent only on a strict subspace of $\mathcal{H}$; thus, unless $\zeta$ lies entirely within this subspace, we have $\mathbb{P}\{\theta_{{\mathsf{K}}_{S}}(\zeta)<c<\infty\}\to 0$ as $\mathsf{K}$$\to\infty$.
Even if all the assumptions required for Theorem 4.4 hold, $\widehat{f}_{\kappa}-f$ converges to a Gaussian random element at a rate depending on $\zeta$ and thus it is not generally possible to construct a uniform confidence band of $f$ from Theorem 4.4 (see Remark 4.5). However, it may be possible to construct a local (or locally approximate) confidence band, which is naturally interpreted. We first note that $f(\zeta)$ may be understood as a partial effect on $y_{t}$ of a perturbation $\zeta$ in $x_{t}$, which is often of interest in practice. If $\mathcal{H}_{y}=\mathbb{R}$ and hence $y_{t}$ is real-valued, $f(\zeta)$ is a real-valued effect on $y_{t}$ of a perturbation $\zeta$, and in this case we may directly use (22) for statistical inference by replacing $\widetilde{C}_{u}$ and $\theta_{{\mathsf{K}}_{S}}$ with their sample counterparts (see Corollary C.1 of the Supplementary Material). Now suppose that $\mathcal{H}_{y}=\mathcal{H}$. In this case, $f(\zeta)$ is a function defined on $a_{1},a_{2}]$, and we may construct a sequence of confidence intervals for local averages of $f(\zeta)$. Specifically, let $\mathcal{I}_{j}=(b_{j+1}-b_{j})^{-1}1\{u\in[b_{j},b_{j+1}]\}$ for some $b_{j}$ and $b_{j+1}$ with $a_{1}\leq b_{j}<b_{j+1}\leq a_{2}$. Then $\langle f(\zeta),\mathcal{I}_{j}\rangle=(b_{j+1}-b_{j})^{-1}\int_{b_{j}}^{b_{j+1}}f(\zeta)(s)ds$ computes the local average of $f(\zeta)$ on the interval $b_{j},b_{j+1}]$. Using the results given in Theorem 4.4, we know that
$\sqrt{T/\theta_{{\mathsf{K}}_{S}}(\zeta)}(\langle(\widehat{f}_{\kappa}(\zeta)-f(\zeta)),\mathcal{I}_{j}\rangle\to_{d}N(0,\langle\mathcal{I}_{j},\widetilde{C}_{u}\mathcal{I}_{j}\rangle).$ (23)
Note that $\langle\mathcal{I}_{j},\widetilde{C}_{u}\mathcal{I}_{j}\rangle$ and $\theta_{{\mathsf{K}}_{S}}$ can also be replaced by their sample counterparts (Corollary C.1 of the Supplementary Material), enabling construction of an asymptotically valid confidence band for the local average via (23). This can be applied to overlapping or non-overlapping sequences of intervals $\{\mathcal{I}_{j}\}_{j=1}^{M}$ with $\cup_{j=1}^{M}\mathcal{I}_{j}=[a_{1},a_{2}]$. These confidence intervals are readily interpretable and can be useful in practice.
@@ 4.4 The model with an intercept
In the previous sections, we developed statistical inferential methods for the case where $\mathbb{E}[y_{t}]=\mathbb{E}[x_{t}]=0$ for simplicity. However, in practice, a nonstationary time series may include a nonzero intercept, and hence our observations $x_{t}$ and $y_{t}$ may satisfy $\mathbb{E}[x_{t}]=\mu_{x}$ and $\mathbb{E}[y_{t}]=\mu_{y}$ for some unknown $\mu_{x}$ and $\mu_{y}$. To accommodate this scenario, one may consider the model with a deterministic term as follows:
$y_{t}=\mu+f(\tilde{x}_{t})+\tilde{u}_{t},$ (24)
where $\mu=f(\mu_{x})-\mu_{y}\in\mathcal{H}$. With a straightforward modification, we can still achieve consistent estimation of $f$ and extend the statistical inference on $f(\zeta)$ given in Section 4.3. More specifically, inference for this case can be implemented using the centered (demeaned) variables ${y}_{c,t}=y_{t}-\bar{y}_{T}$ and $\tilde{x}_{c,t}=x_{t}+e_{t}-\bar{x}_{T}-\bar{e}_{T}$, where $\bar{y}_{T}=T^{-1}\sum_{t=1}^{T}y_{t}$, and $\bar{x}_{T}$ and $\bar{e}_{T}$ are similarly defined. The proposed estimator is given as follows: $\widehat{f}_{c,\kappa}=\widehat{f}_{c,\kappa}^{N}+\widehat{f}_{c,\kappa}^{S},$ where
@@
where $\theta_{c,{\mathsf{K}}_{S}}$ is defined similarly to $\theta_{{\mathsf{K}}_{S}}$ in (D.3), but with $C_{\kappa}^{S}$ and the other operators ($\widetilde{C}_{0}^{S}$, $D_{\kappa}^{S}$, $(D_{\kappa}^{S})_{{\mathsf{K}}_{S}}^{-1}$) depending on $C_{\kappa}^{S}$ being computed from the centered variables ${\rm P}^{S}\tilde{x}_{t}-\mathbb{E}[{\rm P}^{S}\tilde{x}_{t}]$ ($t=1,\ldots,T$) rather than ${\rm P}^{S}\tilde{x}_{t}$. As in the previous case, $\theta_{c,{\mathsf{K}}_{S}}$ and $\widetilde{C}_{u}$ can be replaced by their sample counterparts without affecting the asymptotic result in (25), enabling statistical inference on $f(\zeta)$ in practice. A more detailed discussion including theoretical justification of these results is given in Section C.1.2 of the Supplementary Material.
@@ 5.1 Monte Carlo simulation
We conducted simulation experiments to compare the autocovariance-based estimator with the standard covariance-based one. We compare estimator accuracy and coverage probabilities of confidence intervals for $\langle f(\zeta),\varphi\rangle$ for $\zeta\in\mathcal{H}$ and $\varphi\in\mathcal{H}_{y}$ across estimators. The simulation results support using the autocovariance-based approach with error-contaminated data. However, we postpone these details to Section B of the Supplementary Material to focus on real-data analysis of the economic impact of climate change, which more effectively illustrates our methods’ empirical relevance and usefulness.